The length of a rectangle is 3 m more than twice its width. If the perimeter of the rectangle is 36 cm, find its length and width.
Answer by TimothyLamb(4379)

---
x = 2y + 3
2x + 2y = 36
---
put the system of linear equations into standard form
---
x - 2y = 3
2x + 2y = 36
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= L = 13 m
y= w = 5 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php